Calculate the molality of each of the following solutions:

(a) 15.5 g of sucrose (C12H22O11) in 745 g of water,
__ m
Be sure to answer all parts. Calculate the molality of each of the following solutions: (a) 15.5 g of sucrose (C12H2201n) in 745 g of water, (b) 9.29 moles of ethylene glycol (C2H6O2) in 4985 g of water.
(b) 9.29 moles of ethylene glycol (C2H6O2) in 4985 g of water.
__ m
Answer #1

a) molality = (W /MW) (1000 / mass of solvent in g)

W = 15.5 g

MW = 342 g/mol

mass of solvent = 745 g

molality = (15.5 / 342) (1000 / 745)

molality = 0.061 m

b) molality = n / mass of solvent in Kg

n = 9.29 moles

mass of solvent = 4985 g = 4.985 Kg

molality = 9.29 / 4.985

molality = 1.86 m